OpenStudy (leahmarieed):

need answers quick, will fan + medal Ted practices two types of swimming styles for a total of 50 minutes every day. He spends 20 minutes more practicing the breast stroke than the butterfly stroke. Part A: Write a pair of linear equations to show the relationship between the number of minutes Ted practices the butterfly stroke every day (x) and the number of minutes he practices the breast stroke every day (y). (5 points) Part B: How much time does Ted spend on practicing the breast stroke every day? (3 points) Part C: Is it possible for Ted to have spent 45 minutes practicing the

OpenStudy (_citrus_):

To start with we have: X= minutes of butterfly and Y= minutes of breaststroke. We are told that the total time = 50 minutes. We are also told that however many minutes of butterfly Ted does, he does 20 mins more than breast stroke. Part A: Using the first part of info we know that X+Y=50 Using the second piece of info can represent the total amount of breaststroke time with the following equation: Y=X+ 20 So our linear equations are: X + Y = 50 Y = X + 20 Thats part A

OpenStudy (leahmarieed):

alright, thanks !

OpenStudy (_citrus_):

Part B We can clean up our answers from Part A a lil and make it easier to work with if we solve the first equation for Y. If we do that we get: Y= 50-X Y= X+ 20 Common sense tells you that Y=Y so, Y=Y 50-X= X + 20 2x = 30 X= 15 mins Ted practices the butterfly stroke for 15 mins every day We also know that the total time is 50 mins. If he spends 15 mins on butterfly, simple subtraction tells us that he must spend 35 mins on breaststroke You can check it by plugging the values into the original two equations Thats Part B

OpenStudy (leahmarieed):

alright, i see..

OpenStudy (_citrus_):

Part C We know however much time Ted spends on the butterfly he spends 20 mins more on breast stroke. If he spends 45 mins on butterfly then he must spend 45+20= 65 mins on breast stroke. BUT we already know that he only spends 50 mins total. So, No, it is not possible for Ted to spend 45 mins on butterfly everyday.

OpenStudy (_citrus_):

If you have any questions about any of the parts lmk :)

OpenStudy (leahmarieed):

i actually understand it pretty well now! thank you too much ! appreciated !!!

OpenStudy (_citrus_):

No problem :)